# Multidimensional empirical wavelet transform

### v1

@@ 6 Construction of empirical wavelet systems
For the toy image, Figures 4 and 5 compare the preimage of $\Lambda=\mathrm{B}_{2}(0,\frac{1}{2})$ under the diffeomorphism estimate $\breve{\varphi}_{n}$ to the targeted $\Omega_{n}$ and show the symmetric empirical Gabor wavelet coefficients induced by $\widehat{\psi}^{\mathrm{G}}\circ\breve{\varphi}_{n}$, for the Watershed and Voronoi partitions, respectively, and $n\in\{0,\ldots,9\}$. The diffeomorphisms $\varphi_{n}$ are well estimated for both partitions but the more constrained shapes of the Voronoi partition allow for better estimates. The MSE of the reconstruction from the Watershed and Voronoi transforms are, respectively, $6.93\times 10^{-30}$ and $1.65\times 10^{-31}$, confirming the accurate diffeomorphism estimation. In addition, the different components of the toy image are well recovered by the wavelet coefficients associated with the highest Fourier spectrum energies.
@@
Figure 4: Gabor Watershed transform of the toy image. (Top) Sets $\breve{\varphi}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\varphi_{n}$ mapping the Watershed regions $\Omega_{n}$ (with contour in red) to the disk $\Lambda=\mathrm{B}_{2}(0,1/2)$ and (bottom) resulting empirical Gabor wavelet spectra, for the toy image and $n\in\{0,\ldots,9\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
Figure 5: Gabor Voronoi transform of the toy image. (Top) Sets $\breve{\varphi}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\varphi_{n}$ mapping the Voronoi regions $\Omega_{n}$ (with contour in red) to the disk $\Lambda=\mathrm{B}_{2}(0,1/2)$ and (bottom) resulting empirical Gabor wavelet spectra, for the toy image and $n\in\{0,\ldots,9\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
For the Barbara image, Figures 6 and 7 compare the preimage of $\Lambda=\mathrm{B}_{2}(0,\frac{1}{2})$ under $\breve{\varphi}_{n}$ to the targeted $\Omega_{n}$ and show the symmetric empirical Gabor wavelet coefficients induced by $\widehat{\psi}^{\mathrm{G}}\circ\breve{\varphi}_{n}$, for the Watershed and Voronoi partitions, respectively, and $n\in\{0,\ldots,9\}$. The diffeomorphism estimation is much more accurate on the Vornoi partition than on the Watershed partition, in particular for the region $\Omega_{0}$ which is very large for the Watershed partition. Thus, in terms of reconstruction, Watershed and Voronoi partition lead to MSE of $5.02\times 10^{-25}$ and $2.10\times 10^{-30}$, respectively.
@@
Figure 6: Gabor Watershed transform of the Barbara image. (Top) Sets $\breve{\varphi}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\varphi_{n}$ mapping the Watershed regions $\Omega_{n}$ (with contour in red) to the disk $\Lambda=\mathrm{B}_{2}(0,1/2)$ and (bottom) resulting empirical Gabor wavelet spectra, for the Barbara image and $n\in\{0,\ldots,14\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
Figure 7: Gabor Voronoi transform of the Barbara image. (Top) Sets $\breve{\varphi}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\varphi_{n}$ mapping the Voronoi regions $\Omega_{n}$ (with contour in red) to the disk $\Lambda=\mathrm{B}_{2}(0,1/2)$ and (bottom) resulting empirical Gabor wavelet spectra, for the Barbara image and $n\in\{0,\ldots,14\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
The 1D Shannon mother wavelet is a $sinc$ function given in the Fourier domain by, for every $u\in{\mathbb{R}}$,
$\widehat{\psi}^{\mathrm{1D-S}}(u)=\frac{1}{2\pi}e^{-\imath\pi(u+\frac{3}{2})}\mathbb{1}_{\left(-\frac{1}{2},\frac{1}{2}\right)}(u).$
@@
$\widehat{\psi}^{\mathrm{S}}(u)=\widehat{\psi}^{\mathrm{1D-SH}}(u_{1})\,\widehat{\psi}^{\mathrm{1D-SH}}(u_{2}),$
@@
Assume that the boundaries $\partial\Omega_{n}$ have measures zero. Then, the continuous reconstruction is guaranteed for the systems $\{\psi_{n}\}_{n\in\Upsilon}$ and $\{\phi_{n}\}_{n\in\Upsilon}$ induced by $\widehat{\psi}^{\mathrm{S}}\circ\varphi_{n}$, and is given by Equations 8 and 9.
@@
$\sum_{n\in\Upsilon}\left|\widehat{\psi}_{n}(\xi)\right|^{2}=\sum_{n\in\Upsilon}|\mathrm{det}\;J_{\varphi_{n}}(\xi)|\left|\widehat{\psi}(\varphi_{n}(\xi))\right|^{2}=\frac{|\mathrm{det}\;J_{\varphi_{m}}(\xi)|}{4\pi^{2}}.$
Since $\varphi_{n}$ is a diffomorphism, it follows that, for $a.e.\,\xi\in{\mathbb{R}}^{2}$, $|\mathrm{det}\;J_{\varphi_{m}}(\xi)|>0$ and therefore
@@
Then, Theorem 2 gives a dual frame of $\{\phi_{n}\}_{n\in\Upsilon^{+}}$ permitting the reconstruction following Equation 9. ∎
@@
For the toy image, Figures 8 and 9 compare the preimage of $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ under the diffeomorphism estimate $\breve{\varphi}_{n}$ to the targeted $\Omega_{n}$ and show the symmetric empirical Shannon wavelet coefficients induced by $\widehat{\psi}^{\mathrm{S}}\circ\breve{\varphi}_{n}$, for the Watershed and Voronoi partitions, respectively, and $n\in\{0,\ldots,9\}$. Most of the diffeomorphisms $\varphi_{n}$ are well estimated for both partitions except for the region $\Omega_{4}$ of the Watershed partition. Thus, the MSE of the reconstruction from the Watershed and Voronoi transforms are, respectively, $6.24\times 10^{-10}$ and $3.88\times 10^{-32}$, confirming the higher accuracy of the diffeormophism estimation for the Voronoi partition. However, the components of the toy image are better separated by the wavelet coefficients associated to the Watershed partition.
@@
Figure 8: Shannon Watershed transform of the toy image. (Top) Sets $\breve{\varphi}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\varphi_{n}$ mapping the Watershed regions $\Omega_{n}$ (with contour in red) to the square $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ and (bottom) resulting empirical Shannon wavelet spectra, for the toy image and $n\in\{0,\ldots,9\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
Figure 9: Shannon Voronoi transform of the toy image. (Top) Sets $\breve{\varphi}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\varphi_{n}$ mapping the Voronoi regions $\Omega_{n}$ (with contour in red) to the square $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ and (bottom) resulting empirical Shannon wavelet spectra, for the toy image and $n\in\{0,\ldots,9\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
For the Barbara image, Figures 10 and 11 compare the preimage of $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ under $\breve{\varphi}_{n}$ to the targeted $\Omega_{n}$ along with the symmetric empirical Shannon wavelet coefficients induced by $\widehat{\psi}^{\mathrm{S}}\circ\breve{\varphi}_{n}$, for the Watershed and Voronoi partitions, respectively, and $n\in\{0,\ldots,9\}$. The diffeomorphism estimates associated with the Watershed partition show little accuracy for some regions, in particular regions $\Omega_{0}$, $\Omega_{5}$ and $\Omega_{9}$. In contrast, the Voronoi partition allows for better estimates despite some singularity due to the irregularity at the corner of the square $\overline{\Lambda}$. The reconstruction from the Watershed and Voronoi partitions lead to MSE of, respectively, $9.45\times 10^{-5}$ and $2.19\times 10^{-29}$, which quantifies the higher accuracy of diffeomorphism estimation for the Voronoi partition.
@@
Figure 10: Shannon Watershed transform of the Barbara image. (Top) Sets $\breve{\varphi}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\varphi_{n}$ mapping the Watershed regions $\Omega_{n}$ (with contour in red) to the square $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ and (bottom) resulting empirical Shannon wavelet spectra, for the Barbara image and $n\in\{0,\ldots,14\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
Figure 11: Shannon Voronoi transform of the Barbara image. (Top) Sets $\breve{\varphi}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\varphi_{n}$ mapping the Voronoi regions $\Omega_{n}$ (with contour in red) to the square $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ and (bottom) resulting empirical Shannon wavelet spectra, for the Barbara image and $n\in\{0,\ldots,14\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@ 6.3 Discussion
Overall, the Voronoi partition provides regions that are easier to map than the Watershed partition, particularly when the mother wavelet’s Fourier support is a square. However, the Voronoi partition can lead to a less adapted separation of the harmonic modes, implying that different wavelet coefficients can contain information of the same frequency band. In practice, the mapping can be assessed by the reconstruction error of the inverse wavelet transform to select the most appropriate partitioning method.
@@ 7 Conclusions
In this work, we proposed a general formalism to build multidimensional empirical wavelet systems for a large variety of Fourier domain partitions. In addition, the proposed framework takes into account the potential symmetry of the Fourier supports. Moreover, we showed conditions for the existence of continuous and discrete empirical wavelet frames, which in particular allow to guarantee the reconstruction from the wavelet transforms. Specific wavelet systems based on classic mother wavelets have also been developed and have been shown to be frames under mild assumptions on the Fourier supports. In addition, the implementation toolbox of these wavelet systems will be made freely available at the time of publication. Future work will focus on estimating robustly and efficiently the diffeomorphisms involved in the proposed definitions in 2D. Applications will also be considered.

### v2

@@ 6 Construction of empirical wavelet systems
For the toy image, Figures 4 and 5 compare the preimage of $\Lambda=\mathrm{B}_{2}(0,\frac{1}{2})$ under the diffeomorphism estimate $\breve{\gamma}_{n}$ to the targeted $\Omega_{n}$ and show the symmetric empirical Gabor wavelet coefficients induced by $\widehat{\psi}^{\mathrm{G}}\circ\breve{\gamma}_{n}$, for the Watershed and Voronoi partitions, respectively, and $n\in\{0,\ldots,9\}$. The diffeomorphisms $\gamma_{n}$ are well estimated for both partitions but the more constrained shapes of the Voronoi partition allow for better estimates. The MSE of the reconstruction from the Watershed and Voronoi transforms are, respectively, $6.93\times 10^{-30}$ and $1.65\times 10^{-31}$, confirming the accurate diffeomorphism estimation. In addition, the different components of the toy image are well recovered by the wavelet coefficients associated with the highest Fourier spectrum energies.
@@
Figure 4: Gabor Watershed transform of the toy image. (Top) Sets $\breve{\gamma}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\gamma_{n}$ mapping the Watershed regions $\Omega_{n}$ (with contour in red) to the disk $\Lambda=\mathrm{B}_{2}(0,1/2)$ and (bottom) resulting empirical Gabor wavelet transform spectra, for the toy image and $n\in\{0,\ldots,9\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
Figure 5: Gabor Voronoi transform of the toy image. (Top) Sets $\breve{\gamma}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\gamma_{n}$ mapping the Voronoi regions $\Omega_{n}$ (with contour in red) to the disk $\Lambda=\mathrm{B}_{2}(0,1/2)$ and (bottom) resulting empirical Gabor wavelet transform spectra, for the toy image and $n\in\{0,\ldots,9\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
For the Barbara image, Figures 6 and 7 compare the preimage of $\Lambda=\mathrm{B}_{2}(0,\frac{1}{2})$ under $\breve{\gamma}_{n}$ to the targeted $\Omega_{n}$ and show the symmetric empirical Gabor wavelet coefficients induced by $\widehat{\psi}^{\mathrm{G}}\circ\breve{\gamma}_{n}$, for the Watershed and Voronoi partitions, respectively, and $n\in\{0,\ldots,9\}$. The diffeomorphism estimation is much more accurate on the Vornoi partition than on the Watershed partition, in particular for the region $\Omega_{0}$ which is very large for the Watershed partition. Thus, in terms of reconstruction, Watershed and Voronoi partition lead to MSE of $5.02\times 10^{-25}$ and $2.10\times 10^{-30}$, respectively.
@@
Figure 6: Gabor Watershed transform of the Barbara image. (Top) Sets $\breve{\gamma}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\gamma_{n}$ mapping the Watershed regions $\Omega_{n}$ (with contour in red) to the disk $\Lambda=\mathrm{B}_{2}(0,1/2)$ and (bottom) resulting empirical Gabor wavelet transform spectra, for the Barbara image and $n\in\{0,\ldots,14\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
Figure 7: Gabor Voronoi transform of the Barbara image. (Top) Sets $\breve{\gamma}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\gamma_{n}$ mapping the Voronoi regions $\Omega_{n}$ (with contour in red) to the disk $\Lambda=\mathrm{B}_{2}(0,1/2)$ and (bottom) resulting empirical Gabor wavelet transform spectra, for the Barbara image and $n\in\{0,\ldots,14\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
The 1D Shannon scaling function is a $sinc$ function given in the Fourier domain by, for every $u\in{\mathbb{R}}$,
$\widehat{\psi}^{\mathrm{1D-S}}(u)=e^{-\imath\pi(u+\frac{3}{2})}\mathbb{1}_{\left(-\frac{1}{2},\frac{1}{2}\right)}(u).$
@@
$\widehat{\psi}^{\mathrm{S}}(u)=\widehat{\psi}^{\mathrm{1D-S}}(u_{1})\,\widehat{\psi}^{\mathrm{1D-S}}(u_{2}),$
@@
Assume that the boundaries $\partial\Omega_{n}$ have measures zero. Then, the continuous reconstruction is guaranteed for the empirical wavelet systems $\{\psi_{n}\}_{n\in\Upsilon}$ and $\{\chi_{n}\}_{n\in\Upsilon}$ induced by $\widehat{\psi}^{\mathrm{S}}\circ\gamma_{n}$, and is given by Equations 8 and 9.
@@
$\sum_{n\in\Upsilon}\left|\widehat{\psi}_{n}(\xi)\right|^{2}=\sum_{n\in\Upsilon}|\mathrm{det}\;J_{\gamma_{n}}(\xi)|\left|\widehat{\psi}(\gamma_{n}(\xi))\right|^{2}=|\mathrm{det}\;J_{\gamma_{m}}(\xi)|.$
Since $\gamma_{n}$ is a diffeomorphism, it follows that, for $a.e.\,\xi\in{\mathbb{R}}^{2}$, $|\mathrm{det}\;J_{\gamma_{m}}(\xi)|>0$ and therefore
@@
Then, Theorem 2 gives a dual frame of $\{\chi_{n}\}_{n\in\Upsilon^{+}}$ permitting the reconstruction following Equation 9. ∎
@@
For the toy image, Figures 8 and 9 compare the preimage of $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ under the diffeomorphism estimate $\breve{\gamma}_{n}$ to the targeted $\Omega_{n}$ and show the symmetric empirical Shannon wavelet coefficients induced by $\widehat{\psi}^{\mathrm{S}}\circ\breve{\gamma}_{n}$, for the Watershed and Voronoi partitions, respectively, and $n\in\{0,\ldots,9\}$. Most of the diffeomorphisms $\gamma_{n}$ are well estimated for both partitions except for the region $\Omega_{4}$ of the Watershed partition. Thus, the MSE of the reconstruction from the Watershed and Voronoi transforms are, respectively, $6.24\times 10^{-10}$ and $3.88\times 10^{-32}$, confirming the higher accuracy of the diffeormophism estimation for the Voronoi partition. However, the components of the toy image are better separated by the wavelet coefficients associated to the Watershed partition.
@@
Figure 8: Shannon Watershed transform of the toy image. (Top) Sets $\breve{\gamma}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\gamma_{n}$ mapping the Watershed regions $\Omega_{n}$ (with contour in red) to the square $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ and (bottom) resulting empirical Shannon wavelet transform spectra, for the toy image and $n\in\{0,\ldots,9\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
Figure 9: Shannon Voronoi transform of the toy image. (Top) Sets $\breve{\gamma}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\gamma_{n}$ mapping the Voronoi regions $\Omega_{n}$ (with contour in red) to the square $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ and (bottom) resulting empirical Shannon wavelet transform spectra, for the toy image and $n\in\{0,\ldots,9\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
For the Barbara image, Figures 10 and 11 compare the preimage of $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ under $\breve{\gamma}_{n}$ to the targeted $\Omega_{n}$ along with the symmetric empirical Shannon wavelet coefficients induced by $\widehat{\psi}^{\mathrm{S}}\circ\breve{\gamma}_{n}$, for the Watershed and Voronoi partitions, respectively, and $n\in\{0,\ldots,9\}$. The diffeomorphism estimates associated with the Watershed partition show little accuracy for some regions, in particular regions $\Omega_{0}$, $\Omega_{5}$ and $\Omega_{9}$. In contrast, the Voronoi partition allows for better estimates despite some singularity due to the irregularity at the corner of the square $\overline{\Lambda}$. The reconstruction from the Watershed and Voronoi partitions lead to MSE of, respectively, $9.45\times 10^{-5}$ and $2.19\times 10^{-29}$, which quantifies the higher accuracy of diffeomorphism estimation for the Voronoi partition.
@@
Figure 10: Shannon Watershed transform of the Barbara image. (Top) Sets $\breve{\gamma}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\gamma_{n}$ mapping the Watershed regions $\Omega_{n}$ (with contour in red) to the square $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ and (bottom) resulting empirical Shannon wavelet transform spectra, for the Barbara image and $n\in\{0,\ldots,14\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@
Figure 11: Shannon Voronoi transform of the Barbara image. (Top) Sets $\breve{\gamma}_{n}^{-1}(\Lambda)$ (in white) for the diffeomorphisms $\gamma_{n}$ mapping the Voronoi regions $\Omega_{n}$ (with contour in red) to the square $\Lambda=\left(-\frac{1}{2},\frac{1}{2}\right)\times\left(-\frac{1}{2},\frac{1}{2}\right)$ and (bottom) resulting empirical Shannon wavelet transform spectra, for the Barbara image and $n\in\{0,\ldots,14\}$. The Fourier spectrum energies $\mathrm{E}_{n}$ over the regions $\Omega_{n}\cup\Omega_{-n}$ are indicated above the wavelet spectra.
@@ 6.3 Discussion
As illustrated in this section, continuous wavelet frames can easily be built theoretically but the resulting numerical transform significantly relies on the estimation of the mappings $\gamma_{n}$. Overall, the Voronoi partitions provides regions that are easier to map than the Watershed partitions, particularly when the wavelet kernel’s Fourier support is a square. However, the Voronoi partition can lead to a less adapted separation of the harmonic modes, implying that different wavelet coefficients can contain information of the same frequency band. These observations highlight the need for an efficient diffeomorphism estimator robust to sets $\Lambda$ and $\Omega_{n}$ for extensions to higher dimensions and applications. Moreover, a perfect detection and delimitation of harmonic modes is difficult in practice, which suggests to explore several wavelet kernels in applications.
@@ 7 Conclusions
In this work, we proposed a general formalism to build multidimensional empirical wavelet systems for a large variety of (potentially symmetric) Fourier domain partitions based on continuous wavelet kernels. Moreover, we showed conditions for the existence of continuous and discrete empirical wavelet frames, which in particular allow to guarantee the reconstruction from the wavelet transforms. Specific wavelet systems based on classic wavelet kernels have also been developed and have been shown to be frames under mild assumptions on the Fourier supports. In addition, the implementation toolbox of these wavelet systems will be made freely available at the time of publication. Future work will focus on estimating robustly and efficiently the diffeomorphisms involved in the proposed definitions in 2D and 3D. Applications will also be considered.